OpenStudy (anonymous):

The domain of the function f(x) = \frac ( 37 ) (11 x - 34) is all real numbers except fo

OpenStudy (anonymous):

please please help me!

OpenStudy (anonymous):

solve \[11x-34=0\] in two steps to find your answer

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

do you add 34 to the zero side?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so 3 is the answer?

OpenStudy (anonymous):

hmm no

OpenStudy (anonymous):

it is \[11x-34=0\\ 11x=34\\ x=\frac{34}{11}\] which is close to 3 but not 3 for sure

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

yw